# Higher Gaussian maps on the hyperelliptic locus and second fundamental form

### v1

@@ 1. Introduction
It was proven by Wahl that if $C$ sits on a $K3$ surface, than the first Wahl map $\mu_{1}$ is not surjective (see [18], see also [1]). On the other hand, Ciliberto, Harris and Miranda proved that for the general curve of genus $g\geq 10$, $g\neq 11$, $\mu_{1}$ is surjective (see [2], [16] for another proof).
@@ 3. Rank of higher order Gaussian maps on hyperelliptic curves
Notice that if $g$ is odd, $\dim(Ker(\mu_{g-3})=1$, and if $g$ is even $\dim(Ker(\mu_{g-4})=3$.
@@
(25) $\sum\limits_{1\leq 1<j\leq g-1}a_{ij}Q_{ij}=\sum a_{ij}(ff_{i}dz\odot gf_{j}dz-ff_{j}dz,\odot gf_{i}dz),$
@@
$-\left[\sum\limits_{1\leq i<j\leq g-1}a_{ij}\sum\limits_{r=0}^{k+1}\binom{k+1}{r}f^{(k+1-r)}f_{j}^{(r)}\sum\limits_{e=0}^{k+1}\binom{k+1}{e}g^{(k+1-e)}f_{i}^{(e)}\right]dz^{\otimes(2k+4)}$
(27) $=\left[\sum\limits_{h,l=0}^{k+1}\binom{k+1}{h}\binom{k+1}{l}f^{(k+1-h)}g^{(k+1-l)}\sum\limits_{1\leq i<j\leq l-1}a_{ij}(f_{i}^{(h)}f_{j}^{(l)}-f_{j}^{(h)}f_{i}^{(l)})\right]dz^{\otimes(2k+4)}$
@@
Then using 16 it follows that for every $h+l\leq 2k$
@@
$\left[\binom{k+1}{k}fg^{(1)}(\sum\limits_{1\leq i<j\leq l-1}a_{ij}(f_{i}^{(k+1)}f_{j}^{(k)}-f_{j}^{(k+1)}f_{i}^{(k)})\right]dz^{\otimes(2k+4)}+$
$\left[\binom{k+1}{k}f^{(1)}g(\sum\limits_{1\leq i<j\leq l-1}a_{ij}(f_{i}^{(k)}f_{j}^{(k+1)}-f_{j}^{(k)}f_{i}^{(k+1)})\right]dz^{\otimes(2k+4)}.$
@@
$\left[\binom{k+1}{k}(fg^{(1)}-f^{(1)}g)(\sum\limits_{1\leq i<j\leq l-1}a_{ij}(f_{i}^{(k+1)}f_{j}^{(k)}-f_{j}^{(k+1)}f_{i}^{(k)})\right]dz^{\otimes(2k+4)},$
@@
$(k+1)\mu_{1,L}(t\wedge s)\mu_{2k+1,M}(\sum\limits_{1\leq i<j\leq l-1}a_{ij}(\omega_{i}\wedge\omega_{j})),$
@@
The equation of $Ker(\mu_{2})$ in $I_{2}$ have already been described in the proof of [4, Proposition 4.2]. Indeed we have the following
@@
$Q_{ij}$ as in 3.3. Then $Q\in Ker(\mu_{2})$ if and only if for all $33\leq l\leq 2g-$ 3
@@
Now we generalize the approach in [4]. The strategy is to use Lemma 3.5 together the well-known description of a basis of the $H^{0}(C,K_{C})$ for a hyperelliptic curve $C$. More precisely since we want to describe $Ker(\mu_{2k})$, we can argue locally and suppose that on some open set with coordinate $x$, a basis for $H^{0}(C,K_{C})$ is given by
(29) $\{\frac{dx}{y},x\frac{dx}{y},...,x^{g-1}\frac{dx}{y}\}.$
and the $2:1$ morphism to $\mathbb{P}^{1}$ given by $|L|$ is given by $(x,y)\mapsto x$. Hence a local description of a basis for $H^{0}(C,M)$ is given by
(30) $\{x\frac{dx}{y},...,x^{g-1}\frac{dx}{y}\}.$
@@
with $Q_{ij}$ as in 3.3. For any $k\geq 2$, $Q\in Ker(\mu_{2k})$ if and only if
@@
for all $32k-1\leq l\leq 2g-$ 3. Using lemma 3.5 we have that $\mu_{2k}(Q)=0$ if and only if $\mu_{2k-1,L}(\sum\limits_{1\leq i<j\leq g-1}a_{ij}(\omega_{i}\wedge\omega_{j})=0)$, that is if and only if
@@
where we are using that $dx/y$ is a local frame. Observe that (33) is equal to
@@
So we have shown hat for $Q\in Ker\mu_{2k-2}$, $\mu_{2k}(Q)=0$ if and only if for every $32k-1\leq l\leq 2g-$ 3
@@
Observe that $B^{\prime}_{k,l}$ is just the matrix with rows $w_{l}^{(2)},...,w_{l}^{(2k)}$ where we have removed the entries corresponding to $i+j\neq l$. Notice that these entries are all 00 for all the vectors $w_{l}^{(r)}$.
Define $c_{k,l}=min\{n_{l},k\}$ and let $B_{k,l}$ be the minor of order $c_{k,l}$ given by the first $c_{k,l}$ rows and $c_{k,l}$ columns. We are going to prove that it is not zero for all $33\leq l\leq 2g-$ 3 by induction on $k$.
@@
$B_{2,l}=\begin{pmatrix}j-i&j-1-2\\ ij(j-i)&(j-i-2)(i+1)(j-1).\end{pmatrix}$
@@
$det(B_{2,l})=(j-1)(j-1-2)det\begin{pmatrix}1&1\\ ij&(i+1)(j-1).\end{pmatrix}$
@@
$A_{2,l}:=\begin{pmatrix}1&1\\ ij&(i+1)(j-1).\end{pmatrix}.$
@@
$\begin{pmatrix}1&0\\ ij&j-i-1.\end{pmatrix}.$
whose determinant is $j-i-1>0$. Now assume that for every $2\leq k_{0}\leq k$ and for every $33\leq l\leq 2g-$ 3, $B_{k,l}$ is not zero. We want to prove that the same holds for $k+1$. If $n_{l}\leq k$, then $n_{l}<k+1$, $c_{k+1,l}=min\{n_{l},k\}=n_{l}$ and $B_{k+1,l}=B_{k,l}$. In this case, we have nothing to prove. Assume then $n_{l}>k$. Then $c_{k+1,l}=min\{n_{l},k+1\}=k+1$. Then $B_{k+1,l}$ is the following $k+1\times k+1$ matrix
@@
Now let $s_{k}$ be the number of the indexes $l$ such that $c_{k,l}>c_{k-1,l}$. Observe that the condition $c_{k,l}>c_{k-1,l}$ is equivalent to say that $l$-th equation of $Ker(\mu_{2k})$ is independent from the $l$-th equation of $Ker(\mu_{2}),...,Ker(\mu_{2(k-1)})$. Then we have
@@ 4. Schiffer variations and Weierstrass points
Let $C$ be a smooth curve of genus $g$, take a point $p\in C$ and fix a local coordinate $z$ centred in $p$. For $31\leq n\leq 3g-$ 3, we define the $n^{th}$ Schiffer variation at $p$ to be the element $\xi_{p}^{n}\in H^{1}(C,T_{C})\cong H^{0,1}_{\bar{\partial}}(T_{C})$ whose Dolbeault representative is given by ¯$\frac{\bar{\partial}\rho_{p}}{z^{n}}\frac{\partial}{\partial z}$, where $\rho_{p}$ is a bump function in $p$ which is equal to one in a small neighborhood $U$ containing $p$, $\xi_{p}^{n}=[\frac{\bar{\partial}\rho_{p}}{z^{n}}\frac{\partial}{\partial z}]$. Clearly $\xi_{p}^{n}$ depends on the choice of the local coordinate $z$. Take $31\leq n\leq 3g-$ 3. Consider the exact sequence
@@
Lemma 4.2.
There exists a basis $\omega_{1},...,\omega_{g-1}$ of $H^{0}(K_{C}\otimes L^{\vee})$ such that $ord_{p}\omega_{k}=2g-2k-2$, for all $k=1,...,g-1$.
Proof.
Let us set $M:=K_{C}\otimes L^{\vee}$. By Riemann-Roch we have $h^{0}(M)=h^{0}(K_{C}(-2p))=g-1,$ $h^{0}(M(-p))=h^{0}(M(-2p))=g-2$, since $p$ is a Weierstrass point. Analogously, for every $21\leq k\leq g-$ 2, we get $h^{0}(M(-(2k-1)p))=h^{0}(M(-2kp))=g-k-1$. So we choose a basis of $H^{0}(M)$ as follows:
  • $h^{0}(M(-(2g-4)p))=1$ and we take $\omega_{1}\in H^{0}(M(-(2g-4)p))$, $\omega_{1}\neq 0$.
  • $h^{0}(M(-(2g-6)p))=2$ and we take $\omega_{2}\in H^{0}(M(-(2g-6)p))\setminus H^{0}(M(-(2g-4)p))$, so $H^{0}(M(-(2g-6)p))=\langle\omega_{1},\omega_{2}\rangle$.
@@
  • $h^{0}(M(-2p))=g-2$, and we set $H^{0}(M(-2p))=\langle\omega_{1},...,\omega_{g-2}\rangle$.
  • $h^{0}(M)=g-1$; $\omega_{g-1}\in H^{0}(M)\setminus H^{0}(M(-p))$, and we have $H^{0}(M)=\langle\omega_{1},...,\omega_{g-1}\rangle$.
@@ 5. Second fundamental form of the Hyperelliptic Torelli locus
In [9] a relation between higher even Gaussian maps of the canonical bundle on a smooth projective curve of genus $g\geq 4$ and the second fundamental form of the Torelli map is given. This is a generalisation of a result obtained by Colombo, Pirola and Tortora on the second Gaussian map and the second fundamental form in [8] (see also [6, Theorem 2.2]). In fact, in [6, Theorem 2.2]) it is proven that for any $Q\in I_{2}$, we have
@@ 6. Isotropic subspaces
In this section, given a hyperelliptic curve $C$ of genus $g\geq 3$ and a Weierstrass point $p\in C$, we will determine for any $k\leq\lfloor\frac{g-3}{2}\rfloor$, the maximum $m\leq\lfloor\frac{g-1}{2}\rfloor$, such that the subspace $V_{m}=\langle\xi_{p}^{1},\xi_{p}^{2},...,\xi_{p}^{2m+1}\rangle\subset V_{\lfloor\frac{g-1}{2}\rfloor}\subset H^{1}(T_{C})^{+}$ is isotropic with respect to all the quadrics $\rho(Q)$, $\forall Q\in Ker(\mu_{2k})$. We will show that $m=k$, namely that $V_{k}$ is isotropic for $\rho(Q)$, $\forall Q\in Ker(\mu_{2k})$, while $V_{k+1}$ is not (see Theorems 6.5, 6.7).
@@
$\sum_{\alpha,\beta}b_{\alpha\beta}g_{\alpha}^{(h)}(0)g_{\beta}^{(l)}(0)=0,\ \forall h,l\geq 0,\ h+l\leq m,$
@@
$\rho_{HE}(Q)(\xi_{p}^{n}\odot\xi_{p}^{r})=\rho(Q)(\xi^{n}_{p}\odot\xi^{r}_{p})=2\pi i\left(\sum_{k=0}^{n-1}\left(\sum^{u}_{\alpha,\beta=1}b_{\alpha\beta}f^{(2k+1-l)}_{i}(0)f_{j}^{(l)}(0)\right)\frac{(n-l)}{l!(2k+1-l)!}\right).$
@@
for $11\leq i<j\leq g-$ 1, where $L$ is the $g^{1}_{2}$, $H^{0}(L)=\langle s,t\rangle$, and $H^{0}(M)=H^{0}(K_{C}\otimes L^{\vee})=\langle\omega_{1},...,\omega_{g-1}\rangle$, where the sections $\omega_{i}$ are as in Lemma 4.2. Let $Q$ be a quadric in $I_{2}$, then
$Q=\sum\limits_{1\leq i<j\leq g-1}a_{ij}Q_{ij}=\sum\limits_{\alpha,\beta=1}^{u}b_{\alpha\beta}\gamma_{\alpha}\gamma_{\beta},$
@@
$Q\in Ker(\mu_{2k})\Leftrightarrow\sum\limits_{\alpha,\beta=1}^{u}b_{\alpha\beta}g_{\alpha}^{(h)}g_{\beta}^{(l)}\equiv 0\quad\forall h+l\leq 2k+1.$
@@
$\mu_{2k+2}(Q)=\sum\limits_{\alpha,\beta=1}^{u}b_{\alpha\beta}g_{\alpha}^{(h)}g_{\beta}^{(l)}=\mu_{1,L}(s\wedge t)\mu_{2k+1,M}(\sum_{1\leq i<j\leq g-1}a_{ij}\omega_{i}\wedge\omega_{j})=$
$=(s^{\prime}t-t^{\prime}s)\sum\limits_{1\leq i<j\leq g-1}a_{ij}(\omega_{i}^{(r)}\omega_{j}^{(s)}-\omega_{i}^{(s)}\omega_{j}^{(r))})(dz)^{2k+4}\quad\forall h+l=2k+2,\quad\forall r+s=2k+1.$
@@
$\sum\limits_{\alpha,\beta=1}^{u}b_{\alpha\beta}g_{\alpha}^{(h)}g_{\beta}^{(l)}(p)=0\quad\forall h+l=2k+2,$
@@
$\sum\limits_{\alpha,\beta=1}^{u}b_{\alpha\beta}g_{\alpha}^{(h)}g_{\beta}^{(l)}(p)\neq 0,$
@@
Let $Q=\sum\limits_{1\leq i<j\leq g-1}a_{ij}Q_{ij}\in I_{2}$. If $Q\in Ker(\mu_{2k})$, setting $j=g-s$, we have
$a_{i,g-s}=0\quad\forall g-(2k+2-s)\leq i<g-s$
@@
Recall that in Lemma 3.7 we have shown that $Q\in Ker(\mu_{2k})$, $k\geq 2$, if and only if for every $32\leq l\leq 2g-$ 3,
and for every $m,l$ such that $2\leq m\leq k$, $32m-1\leq l\leq 2g-$ 3,
$\sum\limits_{\begin{subarray}{c}1\leq i<j\\ i+j=l\\ i\geq m-1,\\ j\geq m-1\end{subarray}}a_{ij}(j-i)ij(i-1)(j-1)...(i-(m-2))(j-(m-2))=0.$
As in the proof of Theorem 3.1), for any $33\leq l\leq 2g-$ 3 set $n_{l}=\#\{(i,j):1\leq i<j\leq g-1,\ i+j=l\}$. In Theorem 3.1 we have shown that when $n_{l}\geq k$ the equations of $\mu_{2}$,…$\mu_{2k}$ relative to the index $l$ are linearly independent. When $l=2g-(2k+1)$, or $l=2g-(2k+1)-1$, we have that $n_{l}=k$. This indeed follows from the last part of the proof of Theorem 3.1. Hence when $l$ is equal to $2g-(2k+1)$, or $12g-(2k+1)-$ 1, we have a system of $k$ linearly independent equations in $k$ variables. Applying the same argument for $\mu_{2},...,\mu_{2k}$, this gives $a_{i,g-s}=0$ for $l=i+g-s\geq 2g-(2k+1)-1$, and hence for $i\geq g-(2k+2-s)$ (and $j=g-s>i$).

### v2

@@ 1. Introduction
It was proven by Wahl that if $C$ sits on a $K3$ surface, then the first Wahl map $\mu_{1}$ is not surjective (see [18], see also [1]). On the other hand, Ciliberto, Harris and Miranda proved that for the general curve of genus $g\geq 10$, $g\neq 11$, $\mu_{1}$ is surjective (see [2], [16] for another proof).
@@ 3. Rank of higher order Gaussian maps on hyperelliptic curves
Notice that if $g$ is odd, $\dim(Ker(\mu_{g-3}))=1$, and if $g$ is even $\dim(Ker(\mu_{g-4}))=3$.
@@
(25) $\sum\limits_{1\leq 1<j\leq g-1}a_{ij}Q_{ij}=\sum a_{ij}(ff_{i}dz\odot gf_{j}dz-ff_{j}dz\odot gf_{i}dz),$
@@
$-\left[\sum\limits_{1\leq i<j\leq g-1}a_{ij}\sum\limits_{r=0}^{k+1}\binom{k+1}{h}f^{(k+1-h)}f_{j}^{(h)}\sum\limits_{e=0}^{k+1}\binom{k+1}{l}g^{(k+1-l)}f_{i}^{(l)}\right]dz^{\otimes(2k+4)}$
(28) $=\left[\sum\limits_{h,l=0}^{k+1}\binom{k+1}{h}\binom{k+1}{l}f^{(k+1-h)}g^{(k+1-l)}\sum\limits_{1\leq i<j\leq g-1}a_{ij}(f_{i}^{(h)}f_{j}^{(l)}-f_{j}^{(h)}f_{i}^{(l)})\right]dz^{\otimes(2k+4)}$
@@
Then using (16) and (3) it follows that for every $h+l\leq 2k$
@@
$\left[\binom{k+1}{k}fg^{(1)}(\sum\limits_{1\leq i<j\leq g-1}a_{ij}(f_{i}^{(k+1)}f_{j}^{(k)}-f_{j}^{(k+1)}f_{i}^{(k)})\right]dz^{\otimes(2k+4)}+$
$\left[\binom{k+1}{k}f^{(1)}g(\sum\limits_{1\leq i<j\leq g-1}a_{ij}(f_{i}^{(k)}f_{j}^{(k+1)}-f_{j}^{(k)}f_{i}^{(k+1)})\right]dz^{\otimes(2k+4)}.$
@@
$\left[\binom{k+1}{k}(fg^{(1)}-f^{(1)}g)(\sum\limits_{1\leq i<j\leq g-1}a_{ij}(f_{i}^{(k+1)}f_{j}^{(k)}-f_{j}^{(k+1)}f_{i}^{(k)})\right]dz^{\otimes(2k+4)},$
@@
$(k+1)\mu_{1,L}(t\wedge s)\mu_{2k+1,M}(\sum\limits_{1\leq i<j\leq g-1}a_{ij}(\omega_{i}\wedge\omega_{j})),$
@@
Consider a hyperelliptic curve given by the affine equation:
(29) $y^{2}=\prod_{i=1}^{2g+2}(x-t_{i}).$
For simplicity from now on we will assume $t_{1}=0$, so $|M|=|K_{C}(-2p)|$, where $p$ is the Weierstrass point $(0,0)$ in affine coordinates. Consider the basis
(30) $\{\alpha_{0}=\frac{dx}{y},\alpha_{1}=\frac{xdx}{y},...,\alpha_{g-1}=\frac{x^{g-1}dx}{y}\}$
of $H^{0}(K_{C})$. Then a basis of $H^{0}(M)$ is given by
(31) $\{\alpha_{1}=\frac{xdx}{y},...,\alpha_{g-1}=\frac{x^{g-1}dx}{y}\}.$
The equations of $Ker(\mu_{2})$ in $I_{2}$ in terms of this basis have already been described in the proof of [4, Proposition 4.2]. Indeed we have the following
@@
$Q_{ij}$ as in 3.3, with respect to the basis of $H^{0}(M)$ given by $\alpha_{i}=x^{i}\frac{dx}{y}$, for $i=1,...,g-1$. Then $Q\in Ker(\mu_{2})$ if and only if for all $33\leq l\leq 2g-$ 3
@@
Now we generalize the approach in [4]. The strategy is to use Lemma 3.5 together with the explicit description of the basis $\alpha_{0},...,\alpha_{g-1}$ of the $H^{0}(C,K_{C})$ for a hyperelliptic curve $C$. We have the following result.
@@
with $Q_{ij}=s\alpha_{i}\odot t\alpha_{j}-s\alpha_{j}\odot t\alpha_{i}$ as in 3.3, for $i=1,...,g-1.$ For any $k\geq 2$, $Q\in Ker(\mu_{2k})$ if and only if
@@
for all $32k-1\leq l\leq 2g-$ 3. Using lemma 3.5 we have that $\mu_{2k}(Q)=0$ if and only if $\mu_{2k-1,L}(\sum\limits_{1\leq i<j\leq g-1}a_{ij}(\omega_{i}\wedge\omega_{j}))=0$, that is if and only if
@@
where we are taking $x$ as local coordinate. Observe that (35) is equal to
@@
So we have shown that for $Q\in Ker\mu_{2k-2}$, $\mu_{2k}(Q)=0$ if and only if for every $32k-1\leq l\leq 2g-$ 3
@@
Observe that $B^{\prime}_{k,l}$ is just the matrix with rows $w_{l}^{(2)},...,w_{l}^{(2k)}$ where we have removed the entries corresponding to $i+j\neq l$. Notice that these entries are all 00 for all the vectors $w_{l}^{(2r)}$.
Define $c_{k,l}=min\{n_{l},k\}$ and let $B_{k,l}$ be the minor of $B^{\prime}_{k,l}$ of order $c_{k,l}$ given by the first $c_{k,l}$ rows and $c_{k,l}$ columns. We are going to prove that it is not zero for all $33\leq l\leq 2g-$ 3 by induction on $k$.
@@
$B_{2,l}=\begin{pmatrix}j-i&j-i-2\\ ij(j-i)&(j-i-2)(i+1)(j-1)\end{pmatrix}$
@@
$det(B_{2,l})=(j-i)(j-i-2)det\begin{pmatrix}1&1\\ ij&(i+1)(j-1)\end{pmatrix}$
@@
$A_{2,l}:=\begin{pmatrix}1&1\\ ij&(i+1)(j-1)\end{pmatrix}$
@@
$\begin{pmatrix}1&0\\ ij&j-i-1\end{pmatrix}$
whose determinant is $j-i-1>0$. Now assume that for every $2\leq k_{0}\leq k$ and for every $33\leq l\leq 2g-$ 3, $B_{k,l}$ is not zero. We want to prove that the same holds for $k+1$. If $n_{l}\leq k$, then $n_{l}<k+1$, $c_{k+1,l}=min\{n_{l},k\}=n_{l}$ and $B_{k+1,l}=B_{k,l}$. In this case, we have nothing to prove. Assume then $n_{l}>k$. Then $c_{k+1,l}=min\{n_{l},k+1\}=k+1$. Then $B_{k+1,l}$ is the following $(k+1)\times(k+1)$ matrix
@@
Now let $s_{k}$ be the number of the indexes $l$ such that $c_{k,l}>c_{k-1,l}$. Observe that the condition $c_{k,l}>c_{k-1,l}$ is equivalent to say that $l$-th equations of $Ker(\mu_{2k})$ are independent from the $l$-th equations of $Ker(\mu_{2}),...,Ker(\mu_{2(k-1)})$. Then we have
@@ 4. Schiffer variations and Weierstrass points
Let $C$ be a smooth curve of genus $g>1$, take a point $p\in C$ and fix a local coordinate $z$ centred in $p$. For $31\leq n\leq 3g-$ 3, we define the $n^{th}$ Schiffer variation at $p$ to be the element $\xi_{p}^{n}\in H^{1}(C,T_{C})\cong H^{0,1}_{\bar{\partial}}(T_{C})$ whose Dolbeault representative is given by ¯$\frac{\bar{\partial}\rho_{p}}{z^{n}}\frac{\partial}{\partial z}$, where $\rho_{p}$ is a bump function in $p$ which is equal to one in a small neighborhood $U$ containing $p$, $\xi_{p}^{n}=[\frac{\bar{\partial}\rho_{p}}{z^{n}}\frac{\partial}{\partial z}]$. Clearly $\xi_{p}^{n}$ depends on the choice of the local coordinate $z$. Take $31\leq n\leq 3g-$ 3. Consider the exact sequence
@@
Remark 4.2.
Consider the hyperelliptic curve $C$ with equation (29). Fix the Weierstrass point $p=(0,0)$. For the computations of the next section it is more convenient to use the basis of $H^{0}(M)$ given by $\omega_{i}:=\alpha_{g-i}=\frac{x^{g-i}dx}{y}$, $i=1,...,g-1$. Notice that we have
$ord_{p}\omega_{k}=2g-2k-2,\ for\ all\ k=1,...,g-1.$
Hence we get:
  • $H^{0}(M)=\langle\omega_{1},...,\omega_{g-1}\rangle.$
  • $H^{0}(M(-p))=H^{0}(M(-2p))=\langle\omega_{1},...,\omega_{g-2}\rangle.$
@@
  • $H^{0}(M(-(2g-4)p))=H^{0}(M(-(2g-5)p))=\langle\omega_{1}\rangle.$
So if we take a quadric $Q=\sum\limits_{1\leq i<j\leq g-1}a_{ij}(s\alpha_{i}\odot t\alpha_{j}-s\alpha_{j}\odot t\alpha_{i})\in I_{2},$ as in Lemma 3.6 or Lemma 3.7, we have:
$Q=\sum\limits_{1\leq i<j\leq g-1}a_{ij}(s\alpha_{i}\odot t\alpha_{j}-s\alpha_{j}\odot t\alpha_{i})=-\sum\limits_{1\leq k<h\leq g-1}a_{g-h,g-k}(s\omega_{k}\odot t\omega_{h}-s\omega_{h}\odot t\omega_{k})=$
$=\sum\limits_{1\leq k<h\leq g-1}b_{k,h}(s\omega_{k}\odot t\omega_{h}-s\omega_{h}\odot t\omega_{k}),$
with $b_{k,h}=-a_{g-h,g-k}$.
@@ 5. Second fundamental form of the Hyperelliptic Torelli locus
In [9] a relation between higher even Gaussian maps of the canonical bundle on a smooth projective curve of genus $g\geq 4$ and the second fundamental form of the Torelli map is given. This is a generalisation of a result obtained by Colombo, Pirola and Tortora on the second Gaussian map and the second fundamental form in [8] (see also [6, Theorem 2.2]). In fact, in [6, Theorem 2.2]) it is proven that for any $p\in C$, and for any $Q\in I_{2}$, we have
@@ 6. Isotropic subspaces
In this section, given a hyperelliptic curve $C$ of genus $g\geq 3$ and a Weierstrass point $p\in C$, we will determine for any $k\leq\lfloor\frac{g-3}{2}\rfloor$, the maximum $m\leq\lfloor\frac{g-1}{2}\rfloor$, such that the subspace $V_{m}=\langle\xi_{p}^{1},\xi_{p}^{3},...,\xi_{p}^{2m+1}\rangle\subset V_{\lfloor\frac{g-1}{2}\rfloor}\subset H^{1}(T_{C})^{+}$ is isotropic with respect to all the quadrics $\rho(Q)$, $\forall Q\in Ker(\mu_{2k})$. We will show that $m=k$, namely that $V_{k}$ is isotropic for $\rho(Q)$, $\forall Q\in Ker(\mu_{2k})$, while $V_{k+1}$ is not (see Theorems 6.5, 6.7).
@@
$\sum_{\alpha,\beta}c_{\alpha\beta}g_{\alpha}^{(h)}(0)g_{\beta}^{(l)}(0)=0,\ \forall h,l\geq 0,\ h+l\leq m,$
@@
$\rho_{HE}(Q)(\xi_{p}^{n}\odot\xi_{p}^{r})=\rho(Q)(\xi^{n}_{p}\odot\xi^{r}_{p})=2\pi i\left(\sum_{k=0}^{n-1}\left(\sum^{u}_{\alpha,\beta=1}c_{\alpha\beta}g^{(m+1-k)}_{\alpha}(0)g_{\beta}^{(k)}(0)\right)\frac{(n-k)}{k!(m+1-k)!}\right).$
@@
for $11\leq i<j\leq g-$ 1, where $L$ is the $g^{1}_{2}$, $H^{0}(L)=\langle s,t\rangle$, and $H^{0}(M)=H^{0}(K_{C}\otimes L^{\vee})=\langle\omega_{1},...,\omega_{g-1}\rangle$, where the sections $\omega_{i}$ are as in Remark 4.2. Let $Q$ be a quadric in $I_{2}$, then
$Q=\sum\limits_{1\leq i<j\leq g-1}b_{ij}Q_{ij}=\sum\limits_{\alpha,\beta=1}^{u}c_{\alpha\beta}\gamma_{\alpha}\gamma_{\beta},$
@@
$Q\in Ker(\mu_{2k})\Leftrightarrow\sum\limits_{\alpha,\beta=1}^{u}c_{\alpha\beta}g_{\alpha}^{(h)}g_{\beta}^{(l)}\equiv 0\quad\forall h+l\leq 2k+1.$
@@
$\mu_{2k+2}(Q)=\sum\limits_{\alpha,\beta=1}^{u}c_{\alpha\beta}g_{\alpha}^{(h)}g_{\beta}^{(l)}=\mu_{1,L}(s\wedge t)\mu_{2k+1,M}(\sum_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}\wedge\omega_{j}))=$
$=(s^{\prime}t-t^{\prime}s)\sum\limits_{1\leq i<j\leq g-1}b_{ij}(\omega_{i}^{(r)}\omega_{j}^{(s)}-\omega_{i}^{(s)}\omega_{j}^{(r))})(dz)^{2k+4}\quad\forall h+l=2k+2,\quad\forall r+s=2k+1.$
@@
$\sum\limits_{\alpha,\beta=1}^{u}c_{\alpha\beta}g_{\alpha}^{(h)}g_{\beta}^{(l)}(p)=0\quad\forall h+l=2k+2,$
@@
$\sum\limits_{\alpha,\beta=1}^{u}c_{\alpha\beta}g_{\alpha}^{(h)}g_{\beta}^{(l)}(p)\neq 0,$
@@
Let $Q=\sum\limits_{1\leq r<m\leq g-1}b_{r,m}Q_{r,m}\in I_{2}$, where the quadrics $Q_{r,m}$ are considered with respect to the basis $\{\omega_{i}\}_{i=1,...,g-1}$ given in Remark 4.2. If $Q\in Ker(\mu_{2k})$, we have
$b_{r,m}=0\quad\forall\quad r+m\geq 2g-(2k+2)\qquad(hence\ m\geq g-k).$
@@
Recall that we have $Q=\sum\limits_{1\leq i<j\leq g-1}a_{ij}(s\alpha_{i}\odot t\alpha_{j}-s\alpha_{j}\odot t\alpha_{i})$, where $a_{ij}=-b_{g-j,g-i}$ (see Remark 4.2). We will prove by induction on $k$ that
(47) $a_{ij}=0\quad\forall\quad i+j\leq 2k+2.$
Notice that this finishes the proof. Indeed, putting $r=g-j,m=g-i$ we have
$i+j\leq 2k+2\Leftrightarrow r+m=2g-(i+j)\geq 2g-(2k+2).$
We also get $m\geq g-k$, because if $m\leq g-k-1$, $r<m$ implies $r+m<2g-2k-2$.
Recall that in Lemma 3.6 we have shown that $Q\in Ker(\mu_{2})$ if and only if for every $33\leq l\leq 2g-$ 3,
By Lemma 3.7, $Q\in Ker(\mu_{2k})$, $k\geq 2$, if and only if $Q\in Ker(\mu_{2})$ and and for every $m,l$ such that $2\leq m\leq k$, $32m-1\leq l\leq 2g-$ 3,
So let us prove (47). The case $k=1$ is trivial, since $42k+2=$ 4 and we only have $l=3,4$. From equations (48) we have $a_{12}=a_{13}=0$. Now assume that the result is true for $k-1$, where $k\geq 2$. Hence if $Q\in Ker(\mu_{2(k-1)})$, we have $a_{ij}=0\quad\forall l=i+j\leq 2k$. Now if $Q\in Ker(\mu_{2k})$ then $Q\in Ker(\mu_{2k-2})$, hence $a_{ij}=0\quad\forall l\leq 2k$. So we only have to check the cases $l=2k+1,2k+2$.
  • If $l=2k+1$, then we claim that $i\leq k$. Otherwise, if $i\geq k+1$, we have $j>i\geq k+1$, hence $i+j\geq 2k+3$, a contradiction. In the equations (49) with $l=2k+1$, we have $a_{1,2k},a_{2,2k-1},...,a_{k,k+1}$ as variables, since $i\leq k$. The equations given by (48) and (49) for $m=2,...,k$, $l=2k+1$, are then a system of $k$ equations in $k$ variables, which are linearly independent by the proof of Theorem 3.1. So we get $a_{i,j}=0$ for $i+j=2k+1$.
  • If $l=2k+2$, we get $i\leq k$ exactly as in the previous case. Hence again in (49) with $l=2k+2$, we have exactly $k$ linearly independent equations in the $k$ variables $a_{1,2k+1},a_{2,2k},...,a_{k,k+2}$. So $a_{ij}=0$ if $i+j=2k+2$.